OpenStudy (anonymous):

Solve each system using substitution. Tell whether the system has one solution, infinitely many solutions, or no solution. -x =+ y = -13 3x - y = 19

OpenStudy (marigirl):

can u plz recheck your first equation

OpenStudy (anonymous):

I don't think I typed anything wrong. this is what my question said soooo

OpenStudy (marigirl):

hmm the fist equation says -x =+ y = -13 ..there must only be one equal sign in a equation

OpenStudy (anonymous):

I'm sorry,you're right. take out the equal sign

OpenStudy (marigirl):

which equal sign

OpenStudy (anonymous):

I'll retype the whole thing for you. Solve each system using substitution. Tell whether the system has one solution, infinitely many solutions, or no solution. -x + y = -13 3x - y = 19

OpenStudy (marigirl):

lets make y the subject in the first equation \[y=x-13\] then lets 'substitute' x=13 into the second equation instead of the y \[3x-(x-13)=19\]

OpenStudy (marigirl):

\[3x-x+13=19\]

OpenStudy (anonymous):

so whats the answer? :)

OpenStudy (marigirl):

3x−x+13=19 \[2x=6\] lets solve for x

OpenStudy (anonymous):

get x alone distribute the negative then plug x into the other equation once you solve for y plug into other equation and solve

OpenStudy (marigirl):

so at x=3. this means that at x=3 the graphs intersect

OpenStudy (marigirl):

we can find the y coordinate by plugging in x=3 into any of the two original equations

OpenStudy (marigirl):

What is your final answer (your coordinate) where the graphs intersect?

OpenStudy (marigirl):

and from this u can tell how many times solutions this system of equations have (i.e how many intersections)